How many years will it take for $240000 to amount to $288000 at 8% compound interest per annum?
1
Expert's answer
2011-07-06T06:01:52-0400
The most basic formula to calculate future value is
FV = PV(1 + i)n , where i is the compound interest, n is number of periods.
(1 + i)n = FV/PV
Hence
n = log(1+ i)(FV/PV) = (ln FV - ln PV)/ln (1 + i)
In our case
n = (ln 288,000 - ln 240,000)/ ln (1 + 0.08) = 2.37